OpenStudy (amonoconnor):

"Take the derivative, and solve for x. The equation: y = (x^3) - (26^2) + (160x)" Any and all help is greatly appreciated! :)

OpenStudy (amonoconnor):

I got x = 4, & x = 13.33, but my book has something different...

OpenStudy (shamil98):

Is it asking for dy/dx or d/dx ?

OpenStudy (shamil98):

If we take the derivative with just respect to x, y is treated as a constant.

OpenStudy (amonoconnor):

...it didn't specify, just the generic derivative of the equation, but no u substitution or anything.

OpenStudy (shamil98):

Ok.

OpenStudy (shamil98):

y = (x^3) - (26^2) + (160x) y' = 3x^2 + 160

OpenStudy (amonoconnor):

No -52x ?

OpenStudy (anonymous):

i am assuming that is a typo right?

OpenStudy (shamil98):

you put - (26^2)

OpenStudy (anonymous):

\[y = x^3 - 26x^2+ 160x\]

OpenStudy (shamil98):

if it is a typo then y' = 3x^2 - 52x + 160

OpenStudy (amonoconnor):

Okay, that's what I was getting too.

OpenStudy (amonoconnor):

And x = 4 & 13.33, right??

OpenStudy (amonoconnor):

(Used quadratic formula)

OpenStudy (shamil98):

Set y' = 0 so 0 = 3x^2 - 52x + 160 (x-4)(3x-40) x = 4, x = 40/3 yeah

OpenStudy (amonoconnor):

Alright... Sweet. *God-dang book...

OpenStudy (amonoconnor):

Thank you ! :))

OpenStudy (shamil98):

Yeah, sometimes the book answers are wrong, and you're welcome. :)